### v2

# Unique expansions in number systems via refinement equations

Abstract.
Using the subdivision schemes theory, we develop a criterion to check if any natural number has at most one representation in the $n$-ary number system with a set of non-negative integer digits $A=\{a_{1},a_{2},\ldots,a_{n}\}$ that contains zero. This uniqueness property is shown to be equivalent to a certain restriction on the roots of the trigonometric polynomial $\sum_{k=1}^{n}e^{-2\pi ia_{k}t}$. From this criterion, under a natural condition of irreducibility for $A$, we deduce that in case of prime $n$ the uniqueness holds if and only if the digits of $A$ are distinct modulo $n$, whereas for any composite $n$ we show that the latter condition is not necessary. We also establish the connection of this uniqueness to the semigroup freeness problem for affine integer functions of equal integer slope; this together with the two criteria allows to fill the gap in the work of D.Klarner on Erdös question about densities of affine integer orbits and establish a simple algorithm to check the freeness and the positivity of density when the slope is a prime number.
Key words and phrases:
number system, semigroup freeness, refinement equation, subdivision scheme, Borel measure
1991 Mathematics Subject Classification:
11B75, 20M05, 39A06
1. Number systems and semigroups freeness
Given an integer $n\geq 2$, we consider a set of nonnegative integer numbers $A=\{a_{1},\ldots,a_{m}\}$, which are called digits of the (non-standard) $n$-ary number system $A$. The main results of the paper are obtained assuming $m=n$, but this equality is not necessary in general.
For an integer $k\geq 0$, we call its $A$-expansion any tuple of digits $(\alpha_{0},\ldots,\alpha_{J})$ such that
(1) $k\quad=\quad\sum_{j=0}^{J}\,\alpha_{j}\,n^{j},\,\text{ where }\alpha_{J}\neq 0.$
In particular, for $k=0$ its $A$-expansion is an empty tuple. We say that the $n$-ary number system $A$ has uniqueness property if any integer $k\geq 0$ has at most one $A$-expansion.
The uniqueness property has been studied in the literature mainly because of two applications. The first is the theory of self-similar tiles, where uniqueness gives the existence criterion for the tile. See [14, 20, 21] and references therein for the general multivariate case. The second one is related to the coding theory, but as we shall show, it also turns out to be connected to the freeness of particular semigroups and densities of integer orbit sets arising from the action of integer affine functions.
The active analysis of number systems with positive digits was initiated in early 1980’s by H.A. Maurer, A. Salomaa and D. Wood in [22], where they introduced L-codes and discovered that if the code alphabet is unary, then an L-code corresponds to a non-standard number system [^1] The condition (1) for the most significant digit is set to generalize the definition of uniqueness from [22] to the number systems containing zero, for example the standard digit system $\{0,1,2,\ldots,n-1\}$.. They have proved that such unary code can be uniquely decoded if and only if the corresponding number system has uniqueness property. It was also proved that this property necessarily fails if $m>n$, while if all $\alpha_{i}$ are distinct modulo $n$, then the property holds. The authors also suggested a general decision problem of checking a number system, whether it has the uniqueness property. As it was shown by J. Honkala in [11], this problem is decidable. Some more results concerning number systems were obtained within the next decade; for their description we refer the reader to [12].
Even a little earlier, in [15] D. Klarner came up to a study of another decision problem, which is closely related to the uniqueness of number system. The motivation was coming from Erdös question on when the density of an orbit set $\langle F:S\rangle$ is positive (see [9, p.23] and [19]). Here the set $\langle F:S\rangle$ contains all images of a finite set of non-negative integers $S$ under an application of any number of integer-valued affine functions taken from the set
(2) $F=\{f_{i}=n_{i}x+a_{i},i=1,\ldots,m\},\text{ with $n_{i}\geq~{}2$ and $a_{i}\in{\mathbb{N}}\cup\{0\}$ }.$
Klarner studied a particular case when all slopes $n_{i}$ are equal to some integer $n\geq 2$ and noted that for $m=n$ the (upper) density
$\delta(\langle F:S\rangle)=\lim\sup_{T\to\infty}\frac{\langle F:S\rangle\cup\{0,\ldots T\}}{T+1}$
is equal to 00 if and only if the set $F$ has a non-trivial semigroup relation, i.e.
(3) $f_{u_{1}}\circ\ldots\circ f_{u_{p}}=f_{v_{1}}\circ\ldots\circ f_{v_{q}},$
for some $u_{1},\ldots,u_{p},v_{1},\ldots,v_{q}\in\{1,\ldots,m\}$ such that $(u_{1},\ldots,u_{p})\neq(v_{1},\ldots,v_{s})$.
The decision problem in which, for an input set $F$, one needs to answer whether a relation of the form (3) exists, is known as a freeness problem for a semigroup. [^2] This problem is known to be undecidable for integer upper-triangular $3\times 3$ matrices (see [17] and [1]), and it is still open for integer affine functions, which can be presented by $2\times 2$ upper-triangular matrices.
Klarner noted, that if all slopes in (2) are equal to $n$, the inequality $m>n$ implies that there exists a relation of type (3); for non-equal slopes this was generalized in [16] and [18]. The important subcase $m=n$ was considered separately, but for its resolution the reader was referred to a preprint, which has never appeared. However, it was noted that the freeness can be obtained once the set
$\{0,(a_{2}-a_{1})/d,\ldots,(a_{m}-a_{1})/d\}$
is a complete residue system modulo $n$, where $d=\gcd(a_{2}-a_{1},\ldots,a_{m}-a_{1})$. As we will see in Section 4, in the case of prime $n=m$, the freeness and this condition for residues are actually equivalent.
The uniqueness problem for a number system and the freeness problem for a set of integer-valued affine functions of the same slope can be related by the following observation.
Proposition 1.
Let $F=\{f_{i}(x)=nx+a_{i}\mid i=1,\ldots,m\}$ be the set of affine functions such that $n,a_{2},\ldots,a_{m}\in{\mathbb{N}}$ with $n\geq 2$ and $a_{1}=0$. Then the $n$-ary system $A=\{a_{1},\ldots,a_{m}\}$ has uniqueness property if and only if the set $F$ is a free semigroup basis.
Proof. Suppose first that there is a semigroup relation (3). A straightforward computation shows that $f_{u_{1}}\circ\ldots\circ f_{u_{p}}=n^{p}x+c=f_{v_{1}}\circ\ldots\circ f_{v_{q}}=n^{q}x+d$, where
(4) $c$ $=a_{u_{1}}+a_{u_{2}}n+a_{u_{3}}n^{2}+\ldots+a_{u_{p}}n^{p-1},$
$d$ $=a_{v_{1}}+a_{v_{2}}n+a_{v_{3}}n^{2}+\ldots+a_{v_{q}}n^{q-1}.$
Since $c=d$, from (4) we get two non-equal $A$-expansions of the same number.
Suppose that there exist two tuples $(a_{u_{1}},\ldots a_{u_{p}})$ and $(a_{v_{1}},\ldots a_{v_{q}})$ giving $A$-expansions of the same number $c=d$ in (4). Without loss of generality, we suppose $p\geq q$ and consider $S=f^{(p-q)}_{1}=n^{p-q}x$. Then we immediately obtain, that $L=f_{u_{1}}\circ\ldots\circ f_{u_{p}}=n^{p}+c$ and $R=f_{v_{1}}\circ\ldots\circ f_{v_{q}}\circ S=n^{p}+d$, and $L=R$ is a non-trivial semigroup relation since $u_{p}\neq 0$.
$\Box$
The restriction $a_{1}=0$ in the statement of Proposition 1 is actually somewhat artificial, and it can be bypassed by the following tool:
Proposition 2.
Let $F=\{f_{i}(x)=nx+a_{i}\mid i=1,\ldots,m\}$ be a set of affine functions such that $n,a_{1},a_{2},\ldots,a_{m}\in{\mathbb{Z}}$ and $n\geq 2$. Then for any $s\in{\mathbb{Z}}$ the set $F$ is a free semigroup basis if and only if the set $F=\{\tilde{f}_{i}(x)=nx+(a_{i}-s)\mid i=1,\ldots,m\}$ is a free semigroup basis.
Proof. For the function $g(x)=x-c$, we have $g^{-1}\circ f_{i}\circ g=nx+(a_{i}-cn+c)$. Let us take $c=s/(n-1)$, then $g^{-1}\circ f_{i}\circ g=\tilde{f}_{i}$. Consequently, the existence of a group relation
$f_{u_{1}}\circ\ldots\circ f_{u_{p}}=f_{v_{1}}\circ\ldots\circ f_{v_{q}}.$
is equivalent to the existence of the conjugated relation
$g^{-1}\circ f_{u_{1}}\circ\ldots\circ f_{u_{p}}\circ g=g^{-1}\circ f_{v_{1}}\circ\ldots\circ f_{v_{q}}\circ g,$
which can be also written as
$\tilde{f}_{u_{1}}\circ\ldots\circ\tilde{f}_{u_{p}}=\tilde{f}_{v_{1}}\circ\ldots\circ\tilde{f}_{v_{q}}.$
$\Box$
Propositions 1 and 2 together with Theorem 2 from Section 3 will provide an algorithm to check freeness for the set $F$ in the case $m=n$, which was missing in Klarner’s considerations. If $n$ is prime, the freeness check for the set $F$ (and for the density $\delta(\langle F:S\rangle$) to be positive) can be done much easier by the use of Theorem 4 from Section 4.
2. Subdivision scheme and transition operator
Subdivision schemes are iterative algorithms for linear approximation of functions from their values on a mesh. They originated in late 1980s with N.Dyn, D.Levin, J.Gregory, C.De Boor, S.Dubuc, etc., see [2, 7, 8] for many references. They are widely used now for interpolation and approximation of smooth functions and in modelling of curves and surfaces. Some special cases of subdivisions appeared earlier in works of G. De Rham (cutting angle scheme) [4], G.Chaikin [3], etc. We consider only stationary univariate schemes defined by an integer contraction factor $n\geq 2$ and by a finitely supported sequence of real numbers $\{c_{i}\}_{i\in{\mathbb{Z}}}$. After a suitable shift of numeration it can be assumed that $c_{i}=0$ for all $i\notin\{0,\ldots,N\}$ and $c_{0}c_{N}\neq 0$, when $N\geq 1$. We denote by $l_{\infty}$ the space of bounded sequences $\{x_{k}\}_{k\in{\mathbb{Z}}}$ and by $\delta\in l_{\infty}$ the sequence $(\delta)_{k}=\delta_{0k}$ (the Kronecker symbol).
The subdivision operator $S:l_{\,\infty}\to l_{\,\infty}$ acts on $l_{\,\infty}$ as follows:
(5) $(Sg)_{k}\quad=\quad\sum_{i\in{\mathbb{Z}}}\,c_{k-ni}\,g_{i}$
where $g=(g_{i})_{i\in\mathbb{Z}}\in l_{\infty}$. Furthermore, we consider the transition operator $T$ on the space of compactly supported tempered distributions ${\mathcal{S}}_{0}^{\prime}$:
(6) $\bigl{[}Tf\bigr{]}(t)\quad=\quad\sum_{k\in{\mathbb{Z}}}\,c_{k}\,f(nt-k)\,,\qquad f\,\in\,{\mathcal{S}}_{0}^{\prime}\,.$
There is a simple relation between the subdivision scheme and the transition operator. For every $j\in{\mathbb{N}}$, we have:
(7) $[T^{j}f](t)\quad=\quad\sum_{k\in{\mathbb{Z}}}\ \bigl{(}\,S^{j}\delta\,\bigr{)}_{k}\ f(n^{j}\,t\,-\,k)\,,\qquad f\in{\mathcal{S}}^{\prime}\,.$
The proof can be found, for example, in [2] or verified by a direct computation. The following theorem is well known [24]:
Theorem A. Assume $\sum_{k}c_{k}=n$. Then for every $f\in{\mathcal{S}}_{0}^{\prime}$ such that $(f,{\boldsymbol{1}})=1$, the sequence $T^{j}f$ converges in ${\mathcal{S}}^{\prime}$ to a unique solution $\varphi\in{\mathcal{S}}_{0}^{\prime}$ of the functional equation
(8) $\varphi(t)\quad=\quad\sum_{k\in{\mathbb{Z}}}\,c_{k}\,\varphi(nt-k)\,$
such that $(\varphi,{\boldsymbol{1}})=1$ This solution is supported on the segment $[0,N]$ and satisfies the equation $T\varphi=\varphi$.
Of course, we could have omitted the normalization condition $(f,{\boldsymbol{1}})=1$ and obtain by homogeneity: for every $f\in{\mathcal{S}}_{0}^{\prime}$ the sequence $T^{j}f$ converges to $(f,{\boldsymbol{1}})\,\varphi$. The function $\varphi$ is called in the literature refinable function and equation (8) is called a refinement equation.
To every refinement equation we associate the trigonometric polynomial
[…]
called mask. This is the characteristic function of the sequence $\{c_{k}\}_{k\in{\mathbb{Z}}}$. Computing the Fourier transform of both parts of the refinement equation gives the following equation in the frequency domain:
(9) $\widehat{\varphi}(\xi)\quad=\quad\,{\boldsymbol{c}}(\xi/n)\,\widehat{\varphi}(\xi/n)$
Iterating $j$ times we get
(10) $\widehat{\varphi}(\xi)\quad=\quad\,{\boldsymbol{c}}\,\bigl{(}n^{-1}\xi\bigr{)}\cdots{\boldsymbol{c}}\,\bigl{(}n^{-j}\xi\bigr{)}\,\widehat{\varphi}(n^{-j}\xi).$
Now we focus on the case of nonnegative coefficients $c_{k}$. Some of the results below are known [26], we give their proofs for the convenience of the reader. The following simple observation has been made independently in a number of papers [6, 7, 26, 28].
Proposition 3.
If all $c_{k}$ are nonnegative, then $\varphi$ is a Borel measure of pure type, namely, it is either absolutely continuous (i.e., $\varphi\in L_{1}$) or purely singular.
Proof. Choosing $f\,=\,\chi_{[0,1]}$, we see that the functions $T^{j}f$ are all non-negative and, hence, so is their limit. Thus, $\varphi$ is a nonnegative distribution, that is a Borel measure.
According to the Lebesgue theorem, there exists a unique representation of $\varphi$ in the sum of absolutely continuous and singular measures $\varphi=\varphi_{cont}+\varphi_{sing}$. The uniqueness implies that both $\varphi_{cont}$ and $\varphi_{sing}$ satisfy the refinement equation (8). However, this equation possesses a unique solution up to normalization. Therefore, one of those functions is zero.
$\Box$
Proposition 3 rises the question of separating the cases of absolute continuity and singularity of the measure $\varphi$. The following criterion is proved by applying (10) and the Poisson summation formula:
Proposition 4.
Suppose all $c_{k}$ are nonnegative; then $\varphi\in L_{1}$ if and only if $\widehat{\varphi}(m)=0$ for all integer ${m\neq 0}$. Moreover, in this case $\varphi(t)\leq 1$ almost everywhere.
Proof. (Necessity). Assume that there exists $m\in{\mathbb{Z}}\setminus\{0\}$ such that $\widehat{\varphi}(m)\neq 0$. Taking an arbitrary number $j\in{\mathbb{N}}$ and substituting $\xi=n^{j}m$ to equation (10), we get
$\widehat{\varphi}(n^{j}m)\ =\ {\boldsymbol{c}}\bigl{(}n^{j-1}m\bigr{)}\cdots{\boldsymbol{c}}\bigl{(}m\bigr{)}\widehat{\varphi}\bigl{(}m\bigr{)}\ =\ \widehat{\varphi}\bigl{(}m\bigr{)}.$
[…]
(Sufficiency). If $\widehat{\varphi}(m)=\delta_{\,m0}$, then applying the Poisson summation formula to the function $\varphi(t-\cdot)$ we get
[…]
$\Box$
The criterion of Proposition 4 has a disadvantage that it involves the function $\widehat{\varphi}$, which is a priori unknown. A criterion in terms of the coefficients $\{c_{k}\}_{k\in z}$ exploits the $n$-ary tree. We define the tree ${\mathcal{T}}$ as follows. The root is associated to zero and has $n-1$ children with the numbers $\frac{k}{n},\,k=1,\ldots,n-1$. The further construction is by induction: every vertex $\alpha$, apart from the root, has $n$ children $\frac{\alpha+k}{n},\,k=0,\ldots,n-1$. The root has level zero, its children form the first level, etc. Thus, the $j$th root consists of $(n-1)n^{j-1}$ numbers $\,n^{-j}k\ ,k=1,\ldots,n^{j}-1,\ k\not\equiv 0\,({\rm mod}\,n)$. A subset ${\mathcal{A}}$ of vertices of ${\mathcal{T}}$ is called a minimal cut set if every infinite path along the tree from the root (all paths are without backtracking, the root is not in the path) has exactly one common vertex with ${\mathcal{A}}$. All minimal cut sets are finite. The simplest one is the set of vertices of the first level: ${\mathcal{A}}=\{\frac{1}{n},\ldots,\frac{n-1}{n}\}$.
Proposition 5.
Suppose all $c_{k}$ are nonnegative; then $\varphi\in L_{1}$ if and only if there exists a minimal cut set of the $n$-ary tree ${\mathcal{T}}$ that consists of roots of the mask ${\boldsymbol{c}}$.
Proof. (Necessity). Since $\varphi$ is compactly supported, it follows from the Paley-Wiener theorem that $\widehat{f}$ is an entire function. Therefore, it has finitely many zeros, if any, on the unit disc. This implies that there exists $j\geq 1$ such that $\widehat{\varphi}$ does not have zeros on the $j$th level of ${\mathcal{T}}$, t.e., ${\boldsymbol{c}}(n^{-j}m)\neq 0$ for every natural $m<n^{j},\,m\,\not\equiv 0\,({\rm mod}\,n)$. Let $m=\sum_{k=0}^{j}d_{k}n^{k}\,=\,d_{j}\ldots d_{0}$ be the standard $n$-adic expansion of $m$, possibly, starting with zeros, $d_{0}\neq 0$. Add the digit $d_{j+1}=0$ and substitute $\xi=m$ to equation (10). We obtain
(12) $\widehat{\varphi}(m)\ =\ \,\widehat{\varphi}(n^{-j}m)\,\prod_{q=0}^{j}{\boldsymbol{c}}\,\bigl{(}d_{j+1}\ldots d_{j-q+1}.d_{j-q}\ldots d_{0}\bigr{)}\ =\ \widehat{\varphi}(n^{-j}m)\,\prod_{q=0}^{j}{\boldsymbol{c}}\,\bigl{(}0.d_{j-q}\ldots d_{0}\bigr{)}.$
[…]
(Sufficiency). Assume there exists a minimal cut set ${\mathcal{A}}\subset{\mathcal{T}}$ such that ${\boldsymbol{c}}({\mathcal{A}})=0$. Then taking an arbitrary natural number $m$ and applying formula (12) we obtain $\widehat{\varphi}(m)=0$, which completes the proof for positive $m$. The proof for negative $m$ is the same.
$\Box$
3. Criterion of uniqueness for a number system
Now we consider a set of digits $A=\{a_{1},\ldots,a_{n}\}$ of the (non-standard) $n$-ary number system. We will assume that $0=a_{1}<\ldots<a_{n}$. For any integer $k\geq 0$ we denote by $b(k)$ the total number of its $A$-expansions having form (1) and set formally $b(k)=0$ for any integer $k<0$.
The uniqueness expansion property means that $b(k)\leq 1$ for all $k\in{\mathbb{Z}}$. We are going to see that the sequence $\{b(k)\}_{k\in{\mathbb{Z}}}$ is generated by a subdivision operator $S$ with the following coefficients:
(13) $c_{i}\ =\ \left\{\begin{array}[]{ccl}1&,&i\in A\\ 0&,&{\rm otherwise}\end{array}\right.$
This sequence will be referred to as an indicator sequence of $A$. We start with the following simple observation:
Lemma 1.
The sequence $b(k)$ satisfies the following recurrent relations:
(14) $b\,(nq+d)\ =\ \sum_{\,s\in{\mathbb{Z}}}\,c_{\,ns+d}\ b\,(q-s)\,,\quad d=0,\ldots,n-1,\qquad q\geq 0.$
Proof. In expansion (1) for $k=nq+d$, the digit $\alpha_{0}\in A$ must be such that $\alpha_{0}\equiv d\ ({\rm mod}\,n)$, hence $\alpha_{0}\,=\,ns+d$ for some $s$. Subtracting $d$ from both sides of (1) and dividing by $n$, we get $q-s=\sum_{j}\alpha_{j}\,n^{j-1}$, provided that $ns+d\in A$, i.e., $\alpha_{ns+d}=1$. Thus, the total number of expansions of the number $nq+d$ is equal to the sum of numbers of expansions of $q-s$ over all $s$ such that $\alpha_{ns+d}=1$. This completes the proof.
$\Box$
If we denote $k=nq+d$ and $i=q-s$, then the equation (14) becomes
(15) $b(k)\ =\ \sum_{\,i\in{\mathbb{Z}}}\,c_{\,k-ni}\,b(i).$
The right hand side is precisely the subdivision operator (5). Thus, we obtain
Corollary 1.
Let ${\boldsymbol{b}}^{(j)}\in l_{\infty}$ be the sequence $\{b(k)\}_{k=0}^{n^{j}-1}$, complemented by zeros for $k<1$ and for $k\geq n^{j}$. Then ${\boldsymbol{b}}^{(j)}\,=\,S{\boldsymbol{b}}^{(j-1)}$.
Theorem 1.
For every $j\geq 1$, we have
(16) $b(k)\quad=\quad(S^{j}\delta)_{k}\ ,\qquad k=0,\ldots,n^{j}-1$
where $S$ is the subdivision operator defined by the indicator sequence of $A$.
Thus, the sequence ${\boldsymbol{b}}^{(j)}$ defined in Corollary 1 coincides with the sequence $S^{j}\delta$.
Proof. We argue by induction in $j$. For $j=0$, we have $b(0)=1$ and $(S^{0}\delta)_{0}=(\delta)_{0}=1$, so, this is true. The transfer $j-1\,\mapsto\,j$ is provided by Corollary 1.
$\Box$
Now we formulate the fundamental result. Let $\{c_{k}\}_{k\in{\mathbb{Z}}}$ be the indicator sequence of $A$, ${\boldsymbol{c}}(\cdot)$ be the mask of this sequence, ${\mathcal{T}}$ be the $n$-ary tree.
Theorem 2.
The following assertions are equivalent
a) Every natural number possesses at most one $n$-ary expansion with the digit set $A$;
b) The compactly supported solution of the refinement equation $\varphi(t)=\sum_{i=1}^{n}\varphi(nt-a_{i})$ belongs to $L_{1}({\mathbb{R}})$;
c) The $n$-ary tree ${\mathcal{T}}$ possesses a minimal cut set that consists of roots of the polynomial ${\boldsymbol{c}}(t)=\frac{1}{n}\,\sum_{k=1}^{n}e^{-2\pi ia_{k}t}$.
Proof. The equivalence of b) and c) follows from Proposition 5. Let us prove the equivalence of a) and b).
$\mathbf{a)\,\Rightarrow\,b)}$. Assume the contrary: the property a) holds, i.e., $b(k)\leq 1,\,k\in{\mathbb{Z}}$, but $\varphi\notin L_{1}$. The latter means that $\varphi$ is a purely singular Borel measure (Proposition 3). Consider the transition operator $[Tf](t)\,=\,\sum_{i=1}^{n}\varphi(nt-a_{i})$ and take $f=\chi_{[0,1]}$. By equation (7), we have
$[\,T^{j}f\,]\,(t)\quad=\quad\sum_{k\in{\mathbb{Z}}}\,\bigl{(}S^{j}\delta\,\bigr{)}_{k}\ f(n^{j}t\,-\,k)\quad=\quad\sum_{k\in{\mathbb{Z}}}\,\bigl{(}\,S^{j}\delta\,\bigr{)}_{k}\,\chi_{[n^{-j}k,n^{-j}(k+1)]}\,.$
On the other hand, $T^{j}f\,\to\,\varphi$ due to Theorem A. The convergence is in ${\mathcal{S}}^{\prime}$, which, for nonnegative distributions, means the convergence in measure. Applying now Theorem 1 we obtain $\bigl{(}S^{j}\delta\bigr{)}_{k}=b(k)$ for all $k\leq n^{j}-1$. Since the segments $[n^{-j}k,n^{-j}(k+1)]$ cover the segment $[0,1]$ when $k$ runs from zero to $n^{j}-1$, we conclude that the restriction of the function
$\sum_{k\in{\mathbb{Z}}}\bigl{(}S^{j}\delta\bigr{)}_{k}\,\chi_{[n^{-j}k,n^{-j}(k+1)]}$
to the segment $[0,1]$ is equal to
$f_{j}\ =\ \sum_{k=0}^{n^{j}-1}b(k)\,\chi_{[n^{-j}k,n^{-j}(k+1)]}\,.$
Therefore,
(17) $f_{j}\ =\ T^{j}f\,\bigl{|}_{[0,1]}\ \to\ \varphi\bigl{|}_{[0,1]}\qquad\mbox{{\rm as}}\quad j\to\infty\,.$
By Proposition 3, the function $\varphi$ is of pure type, hence, the assumption $\varphi\notin L_{1}$ implies that $\varphi$ is purely singular. So, its restriction to the segment $[0,1]$ is purely singular as well. On the other hand, if $b(x)\leq 1$ for all $k$, then the function on the right hand side of (17) does not exceed one and hence, so its limit $\varphi\bigl{|}_{[0,1]}$. Thus, the function $\varphi\bigl{|}_{[0,1]}$ is majorized by the Lebesgue measure and is not purely singular.
$\mathbf{b)\,\Rightarrow\,a)}$. Assume that there is a number that has at least two $A$-expansions, the longest of which contains $r$ digits. Let $A_{r}\,=\,\bigl{\{}\sum_{j=0}^{r-1}\alpha_{j}n^{j},\ \alpha_{j}\in A,\,j\leq r\bigr{\}}$. Since $a_{1}=0$, the set $A_{r}$ contains all $A$-presentable numbers of length at most $r$. Hence, at least two elements of $A_{r}$ coincide, so we have $|A_{r}|<n^{r}$. Consider the $r$th power of the transition operator:
(18) $[T^{r}f](t)\quad=\quad\sum_{k\in A_{r}}^{n}b_{r}(k)\varphi(n^{r}t-k)\,,$
where $b_{r}(k)$ is the total number of $A$-expansions of the number $k$ with at most $r$ digits. Clearly, $T^{r}\varphi=\varphi$, hence, the refinement equation with the transition operator $T_{r}$ possesses the same solution $\varphi$. By the Hutchinson theorem [13], there is a unique compact set $Q\subset{\mathbb{R}}$ such that $Q\,=\,\bigcup\limits_{k\in A_{r}}n^{-r}(Q+k)$. This is a fractal generated by affine contractions $n^{-r}(\cdot+k)$. The Lebesgue measure $\mu(Q)$ does not exceed $\sum_{k\in A_{r}}n^{-r}\mu(Q+k)\,=\,|A_{r}|\,n^{-r}\,\mu(Q)$. Since $n^{-r}|A_{r}|<1$, it follows that $\mu(Q)=0$. On the other hand, $T_{r}$ respects the set of Borel measures supported on $Q$. Taking an arbitrary such measure $\psi$ normalized by the condition $(\psi,{\boldsymbol{1}})=1$ and applying Theorem A, we obtain $T^{rj}\psi\,\to\,\varphi$ as $\ j\to\infty$. Therefore, $\varphi$ is also supported on $Q$ and hence, is purely singular.
$\Box$
Remark 1.
Now we see that Theorem 2 together with Proposition 1 and Proposition 2 provide an algorithm to check freeness for a set $F=\{f_{i}=nx+a_{i},i=1,\ldots,n\}$, with $n,a_{1},\ldots,a_{n}\in{\mathbb{Z}}$ and $n\geq 2$. Indeed, the degree of the trigonometric polynomial $\sum_{k=1}^{n}e^{-2\pi ia_{k}t}$ is bounded from above, so is the cardinality of its roots on the segment $[0,1]$. Thus, the possible minimal cut sets of the $n$-ary tree in Theorem 2 also have bounded size, which in turn bounds the height of possible tree vertices in the minimal cut. It follows that the check of all such vertices can be done through their finite enumeration and polynomial evaluation in the corresponding points.
4. Uniqueness of the expansion and digits residues
In this Section we will use the criterion from Theorem 2 to study the uniqueness property for the $n$-ary number system consisting of exactly $n$ digits. Let again $A=\{a_{1},\dots,a_{n}\}$ be the set of nonnegative digits such that $0\leq a_{1}<\dots<a_{n}$. We will say that $n$-ary number system $A$ is irreducible if
(19) $\gcd(a_{2}-a_{1},\dots,a_{n}-a_{1},n)=1.$
Since zero is allowed to be present in the set of digits, we remind that A-expansion of a number $k$ has the form
(20) $k=\sum_{j=0}^{J}\alpha_{j}n^{j},\,\text{ where }a_{J}\neq 0.$
If $k$ has at least one such representation, we will say that $k$ is a representable number.
We will say that $n$-ary number system $A$ has weak uniqueness property if two $A$-expansions of the same length represent different numbers. The following statement is useful for working with sets of digits that contain only positive numbers.
Proposition 6.
Let $A=\{a_{1},\ldots,a_{m}\}$ be a set of digits such that $0<a_{1}<a_{2}<\ldots<a_{m}$. Then the $n$-ary system $A$ has weak uniqueness property if and only if the $n$-ary system $B=\{0,a_{2}-a_{1},\ldots,a_{m}-a_{1}\}$ has uniqueness property.
Proof. We prove both directions by contraposition. First, if $A$ does not have weak uniqueness property, then there exist two tuples $U=(u_{0},u_{1},\ldots,u_{s})$ and $V=(v_{0},v_{1},\ldots,v_{s})$ from $A^{s}$ such that
$\sum_{j=0}^{s}u_{j}n^{j}=\sum_{j=0}^{s}v_{j}n^{j},$
and by subtraction of $a_{0}(1+n+\ldots+a^{s})$ from both sides we obtain
$\sum_{j=0}^{s}(u_{j}-a_{0})n^{j}=\sum_{j=0}^{s}(v_{j}-a_{0})n^{j},$
which gives two distinct tuples $U^{\prime}$ and $V^{\prime}$ from $B^{p}$. By removing most significant digits until they are non-zero, we obtain two tuples refuting the uniqueness property for the system $B$.
Now, suppose that the $n$-ary system $B$ does not have the uniqueness property and there exists $U=(u_{0},u_{1},\ldots,u_{s})\in B^{s}$ and $V=(v_{0},v_{1},\ldots,v_{t})\in B^{t}$ such that
(21) $\sum_{j=0}^{s}u_{j}n^{j}=\sum_{j=0}^{t}v_{j}n^{j}.$
Without loss of generality, we may suppose that $s\geq t$. Consider two tuples $U^{\prime},V^{\prime}\in A^{s+1}$ obtained as $U^{\prime}=(u_{0}+a_{0},u_{1}+a_{0},\ldots,u_{s}+a_{0})$ and $(v_{0}+a_{0},v_{1}+a_{0},\ldots,v_{t}+a_{0},a_{0},\ldots,a_{0})$. They represent the same number in the system $A$, obtained from (21) by adding $a_{0}(1+n+\ldots+a^{s})$.
$\Box$
Now we are ready to prove two results on how the uniqueness property of the $n$-ary system with $n$ digits is related to their residues modulo $n$.
Theorem 3.
Let $n$ be a composite number. Then there exists an irreducible set $A$ consisting of $n$ non-negative digits such that the $n$-ary number system $A$ possesses the uniqueness property, but $A$ contains two numbers congruent modulo $n$.
Proof. Since $n$ is a composite number, there exist integers $n_{1},n_{2}$, both larger than $1$ such that $n=n_{1}n_{2}$. We construct the following set of digits
[…]
As $n_{1}>1$, then by taking $u=0$ we get digits $a_{1}=0$, $a_{2}=1$, so the condition (19) holds. Since $n_{1}$ divides $un_{1}n$ and $0\leq v<n_{1}$ in (22), no digit can be congruent to $n_{1}$ modulo $n$. Hence, the digits do not form the complete system of residues modulo $n$.
From (22) it follows that no digits $a,a^{\prime}$ can satisfy the congruence $a-a^{\prime}\equiv un_{1}\pmod{n}$ for $1\leq u<n_{2}$. Then, considering the difference of two expansions of type (20) modulo $n$ we obtain that if $y,y^{\prime}$ are representable numbers, we also have
(23) $y-y^{\prime}\not\equiv un_{1}\;(\bmod\;n),\,\text{ when }1\leq u<n_{2}.$
Now we will prove that any positive integer has at most one expansion of type (20). Assume the contrary, and let $x$ be the least positive integer having at least two expansions
$x=\sum_{j=0}^{J}\alpha_{j}n^{j}=\sum_{j=0}^{J^{\prime}}\alpha^{\prime}_{j}n^{j}\,\text{ with }\alpha_{J}\neq 0\text{ and }\alpha^{\prime}_{J^{\prime}}\neq 0.$
Then the last digits $\alpha_{0}$ and $\alpha^{\prime}_{0}$ must be distinct: otherwise the number $(x-\alpha_{0})/n$ also has two distinct expansions, but this contradicts the supposition of minimality for chosen $x$. Hence, $\alpha_{0}\neq\alpha^{\prime}_{0}$. Without loss of generality we will presume that $\alpha_{0}<\alpha^{\prime}_{0}$. Then we have
$x=ny+\alpha_{0}=ny^{\prime}+\alpha^{\prime}_{0},$
and $y>y^{\prime}$ are some representable numbers. Thus, $\alpha^{\prime}_{0}-\alpha_{0}=n(y-y^{\prime})$, hence
(24) $\alpha^{\prime}_{0}-\alpha_{0}=0\pmod{n}.$
Then, if the digits $\alpha^{\prime}_{0}$ and $\alpha_{0}$ were formed in $A$ as $\alpha^{\prime}_{0}=u_{1}n_{1}n+v_{1}$ and $\alpha^{\prime}=u_{2}n_{1}n+v_{2}$, from (24) we obtain $v_{1}=v_{2}$, hence $\alpha^{\prime}_{0}-\alpha_{0}=(u_{2}-u_{1})n_{1}n$. Therefore, $y-y^{\prime}=(u_{2}-u_{1})n_{1}$ with $1\leq u_{2}-u_{1}<n_{2}$, which contradicts (23). This proves the uniqueness of the expansion.
$\Box$
Theorem 4.
For an arbitrary prime number $p$ the following holds: an irreducible $p$-ary number system $A$ has a uniqueness property, if and only if the set $A$ does not contain two numbers which are congruent modulo $p$.
Proof. The sufficiency is straightforward (e.g., [22, Theorem 4]), so we prove the necessity.
Let $a_{1}<\dots<a_{p}$ be an irreducible set of non-negative digits such that every number has at most one representation in base $p$ with digits $a_{1},\dots,a_{p}$. We will show that in this case $a_{1},\dots,a_{p}$ form a complete system of residues modulo $p$. To do so, we will use Theorem 2. For any tuple ${\boldsymbol{c}}=(c_{1},\dots,c_{p})\in{\mathbb{Z}}_{\geq 0}^{p}$ we define the polynomial
$P_{\boldsymbol{c}}(z)=\sum_{j=1}^{p}z^{c_{j}}.$
For $m\in{\mathbb{N}}$, a complex number $z$ is a primitive $p^{m}$th root of unity if $z^{p^{m}}=1$, but $z^{p^{m-1}}\neq 1$. Denote ${\boldsymbol{a}}=(a_{1},\dots,a_{p})$ and ${\boldsymbol{b}}=(0,a_{2}-a_{1},\ldots,a_{p}-a_{1})$. Using Proposition 6 we can obtain that the uniqueness property for the $n$-ary system $A$ implies the uniqueness for the system $B=\{0,a_{2}-a_{1},\ldots,a_{p}-a_{1}\}$, to which we can apply Theorem 2. The direction (a)$\Rightarrow$(c) implies that some root of the polynomial $P_{\boldsymbol{b}}$ is a primitive $p^{m}$th root of unity for some $m\in{\mathbb{N}}$, from which it follows that for some $m$ all primitive $p^{m}$th roots of unity are roots of $P_{\boldsymbol{b}}$. Since $P_{\boldsymbol{a}}=e^{-2\pi ia_{0}}P_{{\boldsymbol{b}}}$, the latter is also true for the polynomial $P_{\boldsymbol{a}}$. This will suffice for our purpose.
Recall that $z$ is a $p^{m}$th primitive root of unity if and only if it is a root of the cyclotomic polynomial
$Q(z)=Q_{p^{m}}(z)=\sum_{j=0}^{p-1}z^{jp^{m-1}}.$
Next, we observe that for any $a,\tilde{a}\in{\mathbb{Z}}_{+}$ such that $a\geq\tilde{a}$ and $a\equiv\tilde{a}(\bmod p^{m})$, the polynomial $z^{a}-z^{\tilde{a}}$ is divisible by $z^{p^{m}}-1$ which in turn is divisible by $Q$. Hence, all primitive $p^{m}$th roots of unity are roots of $z^{a}-z^{\tilde{a}}$. For any $j=1,\dots,p$ we take $0\leq\tilde{a}_{j}<p^{m}$ so that $\tilde{a}_{j}\equiv a_{j}(\bmod p^{m})$. Denote
$\tilde{P}(z)=\sum_{j=1}^{p}z^{\tilde{a}_{j}}.$
Then the primitive $p^{m}$th roots of unity are roots of $\tilde{P}$. This implies that $\tilde{P}=QR,$ where $R$ is an integral polynomial of degree $<p^{m-1}$. Moreover, all coefficients of $R$ are nonnegative. We have $\tilde{P}(1)=Q(1)=p$. Hence, $R(1)=1$, and $R(z)=z^{u}$ for some $u$ with $0\leq u<p^{m-1}$.
If $m=1$, then $R(z)=1$, $\tilde{P}(z)=\sum_{j=0}^{p-1}z^{j}$. We see that the multiset $\{\tilde{a}_{1},\dots,\tilde{a}_{p}\}$ is actually the set $\{0,\dots,p-1\}$. This means that $a_{1},\dots,a_{p}$ form the complete system of residues modulo $p$ as desired.
Let $m>1$. Then all numbers $\tilde{a}_{j}$ are congruent to $u$ modulo $p^{m-1}$. Consequently, all numbers $a_{j}$ are also congruent to $u$ modulo $p^{m-1}$. This does not agree with the irreducibility of the set $\{a_{1},\dots,a_{p}\}$.
$\Box$
Remark 2.
If $n$ is prime, Theorem 4 provides an easy freeness test for a set of functions $F=\{f_{i}=nx+a_{i},i=1,\ldots,m\}$, with $n\geq~{}2$ and $a_{i}\in{\mathbb{N}}\cup\{0\}$. First, we increasingly order the set $A=\{a_{1},\ldots,a_{n}\}$ and compute $d=\gcd(a_{2}-a_{1},\ldots,a_{p}-a_{1},p)$. Then, the set $\{0,(a_{2}-a_{1})/d,\ldots,(a_{p}-a_{1})/d\}$ is an irreducible $p$-ary number system, and we are left to check whether its elements are all distinct modulo $p$. Here, the correctness of scaling the coefficients by $d$ can be proved similarly to the correctness of the shift in Proposition 2 by the use of conjugation function $g(x)=dx$.
5. Acknowledgements
The work of the second author and the third author was prepared with the support of the Theoretical Physics and Mathematics Advancement Foundation “BASIS”, grants No. 22-7-1-20-1 and No. 22-7-2-32-1 respectively.